Acube of wood supporting a 200gm mass just floTs in water .when the mass is removed the cube rises by 2cm at equill.fnd side of cube

Grade:11

1 Answers

Arun
6 years ago
Dear Ujjawl
 
Please find the answer to your question
Let the edge of cube be l. When mass is on the cube of wood 
200 g + l3 d wood g = l3 dwater g
⇒ l3 dwood = l3 dwater – 200 … (i) 
When the mass is removed
l3 dwood = (l - 2 ) l2 dwater … (ii) 
From (i) and (ii)
l3 dwater - 200 = (l- 2) l2 dwater 
But dwater = 1
∴ l3 - 200 = l2 (l - 2) 
⇒ l = 10 cm
Regards
Arun(askIITians forum expert)